# MEG: Medical Knowledge-Augmented Large Language Models for Question Answering

### v2

@@ 6 Results
Table 1: Ablation study on the utility of the explicit information encoded in knowledge graph triples. We report accuracy on MedQA. ZS stands for ‘zero-shot’; FT stands for ‘fine-tuning’. Results from Labrak et al. (2024).
MedQA PubMedQA MedMCQA MMLU-Medical Avg MMLU-Medical
Human (pass) 60.0 60.0 Clinical K. Genetics Anatomy P. Medicine C. Biology C. Medicine
Human (expert) 87.0 78.0 90.0
Models based on Llama Models based on Llama
MedAlpaca (7B) 40.1±0.4 73.6±0.3 37.0±0.3 55.1±1.1 51.4 53.1±0.9 58.0±2.2 54.1±1.6 58.8±0.3 58.1±1.3 48.6±0.5
MEDITRON-7B∗‡ 52.0±- 74.4±- 59.2±- 54.2±- 60.0 57.2±- 64.6±- 49.3±- 55.4±- 53.8±- 44.8±-
Meerkat-8B∗‡ 74.2±- - 62.7±- 75.2±- 70.7 74.3±- 76.7±- 74.8±- 75.3±- 76.1±- 74.3±-
MEG-Llama (8B) 66.0±0.2 78.0±0.3 60.6±0.3 74.9±0.7 69.9 72.3±0.5 83.0±1.5 64.5±0.7 79.4±0.3 80.6±0.4 69.4±0.9
Models based on Mistral-Instruct 7B Models based on Mistral-Instruct 7B
Mistral-Instruct-v0.1 42.0±0.2 73.8±${}^{\triangledown}_{\pm{0.4}}$ 46.1±0.1 59.1±1.0 55.3 62.9±0.2 57.0±0.8 55.6±1.0 59.4±0.6 62.5±1.0 57.2±2.1
BioMistral 50.6±0.3 77.5±0.1 48.1±0.2 59.1±1.3 58.8 59.9±1.2 64.0±1.6 56.5±1.8 60.4±0.5 59.0±1.5 54.7±1.0
BioMistral DARE 51.1±0.3 77.7±0.1 48.7±0.1 61.9±1.2 59.9 62.3±1.3 67.0±1.6 55.8±0.9 61.4±0.3 66.9±2.3 58.0±0.5
Meerkat-7B 70.3±- - 60.6±- 70.5±- 67.1 71.6±- 74.8±- 63.2±- 77.3±- 70.8±- 65.2±-
MEG-Mistral1 54.6±0.2 74.6±0.6 56.4±0.4 60.3±0.9 61.5 58.1±0.8 68.7±0.2 54.4±0.5 62.9±0.9 61.1±2.2 56.6±1.0
MEG-Mistral3 60.8±0.2 74.4±0.5 58.4±0.6 68.2±0.4 65.5 64.9±0.2 69.6±0.8 63.0±1.0 72.8±0.4 73.6±0.0 65.2±0.2
Table 2: Main results on four medical multiple-choice question answering benchmarks (left) and fine-grained results on MML-Medical (right). We report accuracy ($\uparrow$) and standard deviation ($\downarrow$), when available, of other 7B and 8B medical open-source models. Avg stands for average across datasets. Results reproduced by Labrak et al. (2024). Reproduced by us with the same data splits used in this work. Results from the original papers.
@@
To fully exploit the KG’s rich structural information, we replace the text triples by node embeddings. This approach also compacts the KG’s information into a much shorter input sequence. Table 2 shows performance on the four medical multiple-choice question answering datasets (left), and a fine-grained evaluation on MMLU-Medical subjects (right). Across all datasets, MEG-Mistral consistently outperforms the baselines’ accuracy, and the merged model DARE, the best BioMistral model in Labrak et al. (2024). The only exception is PubMedQA, where BioMistral surpasses MEG. Their higher accuracy may result from using the artificially labeled training set. Instead, we rather train with the small subset of manually labeled samples to avoid biasing the model towards yes/no answers (see § 5). The high accuracy on MMLU-Medical indicates that MEG retains good generalization capabilities. Current SOTA for 7B models, Meerkat-7B, proves the effectiveness of training on chain-of-thought (CoT) in-domain synthetic data. Future work includes exploring CoT instruction tuning in our phase I of training, exploiting information from graph triples instead of relying only on entity labels.
@@ 6.1 Ablation study
In this section, we evaluate how the choice of graph encoder and mapping network architecture impact MEG-Mistral1’s performance on a downstream task (case study on MedQA).
@@
Table 3: Comparison of MEG-Mistral1’s accuracy on MedQA with different mapping networks. ZS: zero-shot; FT: fine-tuned.
@@ 6.2 Qualitative Analysis
We informally probe MEG to estimate its acquired medical knowledge. We handcraft a set of open-ended questions to query Mistral-Instruct before and after the phase I training on UMLS (§ 4.2), and MEG. As an example, we show an excerpt (irrelevant information to the question is omitted for brevity) of models’ answers to the question ‘What are the active and inactive compounds present in Atorvastatin?’ [^16] This question targets Atorvastatin (C0286650 in UMLS). and highlight the key differences. MEG uses the KGE for Atorvastatin and provides a more knowledgeable response.
@@
7 Analysis and Discussion
Our approach is efficient not only because it trains a small fraction of the LLM’s parameters but also because it seamlessly handles out-of-vocabulary terms. A new term’s KGE can be initialized, for instance, by averaging its one-hop neighbors in the KG, making MEG both lightweight and adaptable to new vocabulary. The efficacy of this method should be evaluated in future work.
@@ 8 Conclusion
We introduce MEG, a novel medical knowledge-augmented LLM based on KGEs for question answering tasks. To the best of our knowledge, we are the first to inject pretrained KGEs into an LLM via a lightweight mapping network, enabling the model to interpret structural graph information from the medical domain. We present a comprehensive evaluation on four medical multiple-choice question benchmarks, revealing that LLMs can highly benefit from the factual information encoded in KG embeddings. Our results suggest that integrating KGEs with LLMs offers a promising path towards specialized language models.
Acknowledgments
We are grateful to CoAStaL members for their comments on earlier versions of this project. Special thanks to Rita Ramos, Constanza Fierro and Jonas F. Lotz for their insightful feedback and to Nicolas Garneau for his support during initial experiments. Laura Cabello was supported by the Novo Nordisk Foundation (grant NNF 20SA0066568). The work of Carlos Bobed was supported by the I+D+i project PID2020-113903RB-I00 (funded by MCIN/AEI/ 10.13039/501100011033) and the project T42_23R (funded by Gobierno de Aragón, Spain).

### v3

@@ 5 Experimental Details
Table 1: Ablation study on the utility of the information encoded in knowledge graph triples expressed in natural language. We report accuracy on MedQA. ZS stands for ‘zero-shot’; FT stands for ‘fine-tuning’. Results from Labrak et al. (2024).
6 Results
We evaluate accuracy on four medical multiple-choice question datasets in three variants of MEG: MEG-Mistral1 and MEG-Mistral3, based on Mistral-7B-Instruct-v0.1 and -v0.3, respectively; and MEG-Llama, based on Llama 3 Instruct (8B). We report average accuracy and standard deviation across three random seeds. Our results in Tables 1 and 2 reveal consistent average improvement across datasets compared to baselines. We adopted both LLMs to show that the benefits of our proposal come regardless the underlying model we took as starting point.
In-prompt graph triples provide useful information
MedQA PubMedQA MedMCQA MMLU-Medical Avg
Models based on Llama
Llama 3 Instruct (8B) 60.3±1.4 75.8±1.4 56.4±0.8 72.8±3.3 66.3
MedAlpaca (7B) 40.1±0.4 73.6±0.3 37.0±0.3 55.1±1.1 51.4
MEDITRON (7B) 52.0±- 74.4±- 59.2±- 54.2±- 60.0
Meerkat-8B 74.2±- - 62.7±- 75.2±- 70.7
RAG-MEDITRON (70B) 49.6±1.4 56.4±2.2 52.7±0.8 65.4±1.4 56.0
RAG-Llama 2 (70B) 44.9±1.4 50.4±2.2 43.1±0.8 54.6±1.5 48.3
MEG-Llama (8B) 66.0±0.2 78.0±0.3 60.6±0.3 74.9±0.7 69.9
Models based on Mistral 7B
Mistral-Instruct-v0.1 42.0±0.2 73.8±${}^{\triangledown}_{\pm{0.4}}$ 46.1±0.1 59.1±1.0 55.3
BioMistral 50.6±0.3 77.5±0.1 48.1±0.2 59.1±1.3 58.8
BioMistral DARE 51.1±0.3 77.7±0.1 48.7±0.1 61.9±1.2 59.9
Meerkat-7B 70.3±- - 60.6±- 70.5±- 67.1
RAG-Mixtral (8$\times$7B) 60.0±1.4 67.6±2.1 56.4±0.8 75.9±1.3 65.0
MEG-Mistral1 54.6±0.2 74.6±0.6 56.4±0.4 60.3±0.9 61.5
MEG-Mistral3 60.8±0.2 74.4±0.5 58.4±0.6 68.2±0.4 65.5
Table 2: Main results on four medical multiple-choice question answering benchmarks. We report accuracy ($\uparrow$) and standard deviation ($\downarrow$), when available, of other 7B and 8B medical open-source models. Avg stands for average across datasets. Llama 3 Instruct and Mistral-Instruct are our base, general purpose models. Models with gray background are baselines fine-tuned on biomedical data. MEG variants outperformed the respective baselines. Results reproduced by Labrak et al. (2024). Results from Xiong et al. (2024). Reproduced by us with the same data splits used in this work. Results from the original papers.
@@ 6 Results
To fully exploit the KG’s rich structural information, we use node embeddings to guide the LLM’s generation. This approach also compacts the KG’s information into a much shorter input sequence. Table 2 reports accuracy on four multiple-choice question datasets. Across all datasets, MEG-Llama and both MEG-Mistral consistently outperform the base LLMs and the fine-tuned baselines (in gray). The only exception is PubMedQA, where BioMistral surpasses both MEG-Mistral. Its higher accuracy may result from using the artificially labeled training set. Instead, we rather train with the small subset of manually labeled samples to avoid biasing the model towards yes/no answers (see § 5). The high accuracy on MMLU-Medical indicates that MEG retains good generalization capabilities. See Appendix B for a fine-grained evaluation on MMLU-Medical subjects. Current SOTA for 7B models, Meerkat-7B, proves the effectiveness of training on chain-of-thought (CoT) in-domain synthetic data. Future work includes exploring CoT instruction tuning in our phase I of training, exploiting information from graph triples instead of relying only on entity labels. For reference, we also include RAG results from Xiong et al. (2024). Despite models used by Xiong et al. (2024) are larger than ours and we cannot directly compare to them, accuracy results indicate that MEG is a solid, lightweight alternative.
@@ 6.1 Ablation study
In this section, we take MEG-Mistral1 as reference and evaluate the impact of the choice of graph encoder and mapping network architecture on a downstream task (case study on MedQA).
@@
Table 3: Ablation study on the mapping network done on a separated 1K subset of the train set. Results show accuracy on MedQA for MEG-Mistral1. ZS: zero-shot; FT: fine-tuned.
@@ 6.2 Qualitative Analysis
We informally probe MEG to estimate its acquired medical knowledge. We handcraft a set of open-ended questions to query Mistral-Instruct before and after the phase I training on UMLS (§ 4.2), and MEG-Mistral1. As an example, we show an excerpt (irrelevant information to the question is omitted for brevity) of models’ answers to the question ‘What are the active and inactive compounds present in Atorvastatin?’ [^18] This question targets Atorvastatin (C0286650 in UMLS). and highlight the key differences. MEG uses the KGE for Atorvastatin and provides a more knowledgeable response.
@@
7 Discussion and Future Research Lines
Our approach is efficient not only because it trains a small fraction of the LLM’s parameters but also because it seamlessly handles out-of-vocabulary terms. We present and efficient approach to domain specialization of models, as it updates only a small fraction of the LLM’s parameters while seamlessly handling out-of-vocabulary terms. A new term’s KGE can be initialized, for instance, by averaging its one-hop neighbors in the KG, making MEG both lightweight and adaptable to new vocabulary. However, the efficacy of this method should be evaluated in future work.
@@ 8 Conclusion
We introduce MEG, a novel medical knowledge-augmented LLM based on KGEs for question answering tasks. To the best of our knowledge, we are the first to inject pretrained KGEs into an LLM via a lightweight mapping network. The LLM interprets the graph information encoded in KGEs, which further guides the response generation. We present a comprehensive evaluation on four medical multiple-choice question benchmarks, revealing that LLMs can highly benefit from the factual information encoded in KG embeddings. Our results suggest that integrating KGEs with LLMs offers a promising path towards specialized language models.
Appendix A Training Details
A.1 Phase I: Embedding Transfer Learning
Since the input prompt has a fixed size (see Figure 2), we use a reduced sequence length of 124 to optimize computational efficiency. We train for one epoch with gradient accumulation over 8 steps to achieve an effective batch size of 128. We employ a cosine learning rate scheduler with learning rate of $51e-$ 5, warmup ratio of 3% and no weight decay. We use mixed precision (bfloat16) and FlashAttention2 Dao (2023) to optimize memory usage and speed up computations on the LLM.
A.2 Fine-tuning on Downstream Tasks
Fine-tuning is done for 3 epochs, with a sequence length of 400 (500 for PubMedQA), learning rate of $41e-$ 4 and effective batch size of 32. The remaining hyperparameters have the same value as in Phase I. We did not optimize hyperparameters for Llama-3-Instruct. Total number of trainable parameters: 83M, in both MEG-Mistral and MEG-Llama.
Appendix B Extended Results
Clinical K. Genetics Anatomy P. Medicine C. Biology C. Medicine
Models based on Llama
LLama 3 Instruct (8B) 73.6±2.7 85.0±3.6 68.9±4.0 69.1±2.8 78.5±3.4 61.8±3.7
MedAlpaca (7B) 53.1±0.9 58.0±2.2 54.1±1.6 58.8±0.3 58.1±1.3 48.6±0.5
MEDITRON-7B∗‡ 57.2±- 64.6±- 49.3±- 55.4±- 53.8±- 44.8±-
Meerkat-8B∗‡ 74.3±- 76.7±- 74.8±- 75.3±- 76.1±- 74.3±-
MEG-Llama (8B) 72.3±0.5 83.0±1.5 64.5±0.7 79.4±0.3 80.6±0.4 69.4±0.9
Models based on Mistral 7B
Mistral-Instruct-v0.1 62.9±0.2 57.0±0.8 55.6±1.0 59.4±0.6 62.5±1.0 57.2±2.1
BioMistral 59.9±1.2 64.0±1.6 56.5±1.8 60.4±0.5 59.0±1.5 54.7±1.0
BioMistral DARE 62.3±1.3 67.0±1.6 55.8±0.9 61.4±0.3 66.9±2.3 58.0±0.5
Meerkat-7B 71.6±- 74.8±- 63.2±- 77.3±- 70.8±- 65.2±-
MEG-Mistral1 58.1±0.8 68.7±0.2 54.4±0.5 62.9±0.9 61.1±2.2 56.6±1.0
MEG-Mistral3 64.9±0.2 69.6±0.8 63.0±1.0 72.8±0.4 73.6±0.0 65.2±0.2
Table 4: Performance on six individual MMLU-Medical categories. We report accuracy ($\uparrow$) and standard deviation ($\downarrow$), when available, of other 7B and 8B medical open-source models. Models with gray background are baselines fine-tuned on biomedical data.